### v3

# Atomic Josephson Parametric Amplifier

Abstract
We study the dynamics of a driven atomic Josephson junction that we propose as a parametric amplifier. By periodically modulating the position of the barrier, we induce a small current across the junction, serving as our input signal. The pump field is implemented by modulating the barrier height at twice the Josephson plasma frequency. The resulting dynamics exhibit parametric amplification of the signal through nonlinear mixing between the signal and pump fields, which is encoded in a specific microscopic pattern of density waves and phase excitations that can be addressed within the experimental cold atoms capabilities. This work paves the way for tunable amplifiers in atomtronic circuits, with potential applications in several fields including precision measurements and quantum information processing. At the same time, our analysis provides the microscopic explanation of the general notion of parametric amplification occurring in nonlinear coherent devices.
Parametric amplification is a fundamental concept in nonlinear physics: When the system’s natural frequency $\omega_{0}$ is periodically modulated in time with frequency $\omega_{\mathrm{d}}$ (the pump), a nearly resonant condition can be achieved for $\omega_{\mathrm{d}}\approx 2\omega_{0}$; in this case, an energy transfer from the pump to the system is induced [1]. This idea was implemented for classical electronics [2] and then extended to quantum optics [3, 4], semiconductors [5] and superconducting materials [6, 7, 8, 9, 10]. Parametric amplifiers are extensively used in important technological sectors as radio astronomy [11], communications [12], radar systems [13], spectroscopy [14] and sensing [15]. In quantum technology, parametric amplifiers have become relevant for many applications ranging from the generation of squeezed states [6, 7, 16], entanglement and nonclassical states [17, 18], and quantum-limited amplification [19] to quantum computing [20], quantum sensing and metrology [21, 22].
Superconducting Josephson junctions provide non linear circuits that have been thoroughly employed to achieve a parametric amplification. These devices amplify weak microwave signals with minimal added noise thus generating high-quality amplification. When we apply a strong pump field with frequency $\omega_{\mathrm{d}}$ at twice the Josephson plasma frequency $\omega_{0}$, i.e., $I_{\mathrm{pump}}(t)\sim\sin(2\omega_{0})$, and a weak signal field $I_{\mathrm{signal}}(t)\sim\cos(\omega_{\mathrm{s}}t)$, with frequency $\omega_{\mathrm{s}}\approx\omega_{0}$, the nonlinear inductance of the Josephson junction produces parametric amplification through a nonlinear mixing process between the signal and pump fields [23]. $\omega_{0}$ is determined by $\omega_{0}=\sqrt{2eI_{c}/(\hbar C)}$, where $2e$ is the electronic charge, $I_{c}$ the critical current, and $C$ the capacitance. The idler modes at frequency $\omega_{\mathrm{I}}=\omega_{\mathrm{d}}-\omega_{s}$ ensure the energy conservation. The signal mode at $\omega_{\mathrm{s}}$ is amplified as the pump energy is transferred to it through the nonlinear mixing process.
Ultracold atoms have emerged as an ideal platform for implementing and exploring atomtronic analogs of superconducting circuits [24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. One breakthrough was the realization of Josephson junctions (JJs) using weak links of ultracold atom clouds [34, 35, 36, 37, 38]. This has not only enabled the study of important quantum effects, such as dc-ac Josephson effects [39, 40, 41], current-phase relation [42, 40], Josephson-tunneling enhanced superfluidity [43], with remarkable accuracy and control of the physical conditions, but has also laid the foundation for the development of atom-based devices, such as atomic quantum interference devices (AQUIDs) [44, 28, 29, 45, 30]. Theoretical studies have examined Josephson effects [46, 47, 48], the implementation of universal set of logic gates [49], atomic conductivity in pump-probe setting [50], or driven dynamics [47, 51, 52, 53]. Recently, quantized plateaus of chemical potential were predicted in driven atomic JJs [54], which are the analogue of Shapiro steps in superconducting JJs [55, 56]. Atomic Shapiro steps were measured in Refs. [57, 58], serving as a basis for a pressure standard in atomic condensates.
With remarkable progress in the field of ultracold atoms, driven matter-wave circuits are now well within experimental reach. Here, we propose the implementation of a Josephson parametric amplifier (JPA) using ultracold atoms. Our protocol relies on an atomic JJ in which we periodically modulate both the position and height of the tunnel barrier, see Figs. 1(a, b). This can be realized in ultracold atom experiments using digital micromirror devices [57, 58]. In our system, the modulation of the barrier position serves as the input current signal, while the modulation of the barrier height acts as a pump field. By tuning the pump frequency to twice the Josephson plasma frequency, we induce nonlinear mixing between the pump and signal modes, see Fig. 1(c). A strong pump field drives the system into a regime where parametric amplification occurs. Our results are obtained using classical-field dynamics, which include fluctuating bosonic fields beyond the mean-field description [59, 60, 61, 62]. Additionally, the system dynamics are well captured by a driven circuit model.
Figure 1: Atomic Josephson parametric amplifier (JPA). (a) Simulation of a Josephson junction, which consists of two 2D clouds separated by a Gaussian tunnel barrier of height $V_{0}$ and width $w$ (indicated by two dotted vertical lines). We use $V_{0}/\mu=1.5$ and $w/\xi=1.1$, where $\mu$ is the mean-field energy and $\xi$ is the healing length. (b) Sketch of the JPA protocol. We periodically modulate the barrier position using $x(t)=x_{0}+x_{1}\sin(2\pi f_{\mathrm{s}}t)$, where $x_{1}$ is the amplitude and $f_{\mathrm{s}}$ is the frequency. For the pump we modulate the barrier height, i.e., $V(t)=V_{0}\bigl(1+A_{\mathrm{d}}\cos(2\pi f_{\mathrm{d}}t)\bigr)$, where $A_{\mathrm{d}}$ is the amplitude and $f_{\mathrm{d}}$ is the frequency. (c) Sketch of the operation of a JPA. We modulate the barrier height near twice the plasma frequency $f_{0}$ of the $J$. The signal mode at $f_{\mathrm{d}}/2+\delta f$ is amplified and an idler mode at frequency $f_{\mathrm{d}}/2-\delta f$ is created, with $\delta f$ being the tuning parameter.
Figure 2: JPA dynamics. Time evolution of the imbalance $z(t)$ demonstrating the amplification for the signal at frequency $f_{\mathrm{s}}=30\,\mathrm{Hz}$ and two different sets of signal and pump amplitudes: (a) $v_{1}/v_{c}=0.21$ and $A_{\mathrm{d}}=0.9$. (b) $v_{1}/v_{c}=0.42$ and $A_{\mathrm{d}}=0.5$. $v_{c}$ is the critical velocity above which the junction is resistive. The results at $A_{\mathrm{d}}=0$ represent the system without the pump. $T_{\mathrm{s}}=1/f_{\mathrm{s}}$ is the signal oscillation period. (c, d) Power spectrum of the time evolution in (a, b).
Figure 3: Nonlinear mixing in the power spectrum $S_{\omega}=|z(\omega)|^{2}$ for signal frequencies between $f_{\mathrm{s}}=28$ and $52\,\mathrm{Hz}$ at $\tilde{I}_{1}\equiv I_{1}/I_{c}=0.42$. For each column the spectrum is normalized by its maximum value $S_{\omega,\mathrm{m}}$. (a) Without pump the spectrum mainly displays the central peak at $f_{\mathrm{s}}=f$. (b, c) In the presence of pump there is another peak at $f_{\mathrm{d}}-f_{\mathrm{s}}$ (indicated by white dashed line), located at the frequency of the signal mirrored at the plasma frequency $f_{0}$. The case of $f_{\mathrm{s}}=f_{0}$, where signal and idler are at the same frequency, is a special case known as the degenerate case. (d) The result of the RSJ circuit model for $\tilde{I}_{1}=0.42$ and $A_{\mathrm{d}}=0.8$; see text.
System and protocol.— We consider a homogeneous cloud of bosons confined in a box of dimensions $L_{x}\times L_{y}$. We study the dynamics of this system using classical-field dynamics within the truncated Wigner approximation [59, 60, 61, 62]. The system is described by the Hamiltonian
$\hat{H}_{0}=\int\mathrm{d}{\bf r}\Big[\frac{\hbar^{2}}{2m}\nabla\hat{\psi}^{\dagger}({\bf r})\cdot\nabla\hat{\psi}({\bf r})+\frac{g}{2}\hat{\psi}^{\dagger}({\bf r})\hat{\psi}^{\dagger}({\bf r})\hat{\psi}({\bf r})\hat{\psi}({\bf r})\Big].$ (1)
$\hat{\psi}$ ($\hat{\psi}^{\dagger}$) is the bosonic annihilation (creation) operator. The interaction $g=\tilde{g}\hbar^{2}/m$ is given in terms of the dimensionless parameter $\tilde{g}=\sqrt{8\pi}a_{s}/\ell_{z}$, where $m$ is the mass, $a_{s}$ is the s-wave scattering length and $\ell_{z}=\sqrt{\hbar/(m\omega_{z})}$ is the harmonic oscillator length in the transverse direction. In the limit of a large atom number $N$, we employ the classical-field representation, where the operators $\hat{\psi}$ in Eq. (1) and the equations of motion are replaced by complex numbers $\psi$. The initial states $\psi({\bf r},t=0)$ are generated in a grand canonical ensemble with chemical potential $\mu$ and temperature $T$ via a classical Metropolis algorithm. This distribution captures fluctuations of $\psi({\bf r},t=0)$ around its mean field value. Finally, each initial state is evolved according to the classical equations of motion
$i\hbar\dot{\psi}({\bf r},t)=\Bigl(-\frac{\hbar^{2}}{2m}\nabla^{2}+V({\bf r},t)+g|\psi|^{2}\Bigr)\psi({\bf r},t),$ (2)
[…]
$S_{\omega}=|z(\omega)|^{2},$ (3)
where $z(\omega)=(1/\sqrt{T_{n}})\int dt\,z(t)\exp(-i\omega t)$ is the Fourier transform of $z(t)$. $T_{n}$ is the sampling time for the numerical Fourier transform. The junction displays a well-defined resonance at the Josephson plasma frequency $f_{0}$, which we determine by analyzing $S_{\omega}$ at varying frequency, obtaining $f_{0}=39.4\,\mathrm{Hz}$ in the limit of $v_{1}\rightarrow 0$ (see End Matter for details). For the parametric pump we modulate the barrier height, i.e., $V(t)=V_{0}\bigl(1+A_{\mathrm{d}}\cos(2\pi f_{\mathrm{d}}t)\bigr)$, where $A_{\mathrm{d}}$ is the amplitude and $f_{\mathrm{d}}$ is the frequency, as depicted in Fig. 1(b). This alters the nonlinear inductance of the junction, which gives rise to amplification of the signal due to energy exchange between different modes. When $V(t)$ is modulated at twice the plasma frequency, the system undergoes parametric amplification, where a signal at frequency $f_{\mathrm{d}}/2+\delta f$ is amplified and an idler at frequency $f_{\mathrm{d}}/2-\delta f$ is created, with $\delta f$ being the tuning parameter, see Fig. 1(c). In Fig. 2, we demonstrate the parametric enhancement for the signal at $f_{\mathrm{s}}=30\,\mathrm{Hz}$ with respect to the system without pump. The system quickly reaches the steady state after 232-3 signal cycles and the enhancement persists over long times as required for robust amplification. The power spectrum exemplifies the amplification in frequency space.
Figure 4: JPA characteristics. (a) Dimensionless gain $G_{\mathrm{s}}=S_{\omega}(A_{\mathrm{d}})/S_{\omega}(0)-1$ as a function of $A_{\mathrm{d}}$ is shown for signal amplitudes between $\tilde{I}_{1}=0.12$ and $0.42$. The results of the RSJ model are indicated by the dashed lines. (b) Signal-to-noise ratio (SNR) corresponding to the results shown in (a). (c, d) Simulation and RSJ model results at temperature $T/T_{0}=0.3$, using the same relative signal frequency with respect to the Josephson plasma frequency as in panels (a, b); see text.
Characterization of the JPA dynamics.— To characterize the dynamical regimes we map out the imbalance spectrum for signal frequencies between $f_{\mathrm{s}}=20$ and $100\,\mathrm{Hz}$, which exhibits JPA resonances and other excitation peaks  [^1] See Supplemental Material for the driven junction spectrum, the current-chemical potential characteristic, and the driven RSJ circuit model, which includes Refs. [54, 66].. Fig. 3 highlights the interplay between the signal and idler modes for the lowest JPA resonance. The mode weights are controlled by the pump amplitude. To describe the JPA dynamics we introduce a driven RSJ (resistively shunted junction) circuit model. The barrier velocity results in a modulation of the current across the junction given by $I_{1}\cos(\omega_{s}t)$, where $I_{1}=v_{1}I_{c}/v_{c}$ in which $v_{c}$ is the critical velocity associated with the critical current $I_{c}$ [65]. The barrier height modulation changes the tunneling energy, which can be incorporated as $I_{c}\bigl(1+A_{\mathrm{d}}\sin(\omega_{\mathrm{d}}t)\bigr)$. The Kirchhoff law of the driven RSJ circuit reads
$I_{1}\cos(\omega_{s}t)=I_{c}\bigl(1+A_{\mathrm{d}}\sin(\omega_{\mathrm{d}}t)\bigr)\sin\phi-G\Delta\mu,$ (4)
where $\phi=\phi_{L}-\phi_{R}$ is the phase difference across the junction and $G$ is the conductance [65]. The Josephson relation for the phase dynamics is $\hbar\dot{\phi}=-\Delta\mu$, with $\Delta\mu$ playing the role of the voltage across the junction. We numerically solve Eq. 4 for various signal and pump parameters and obtain the time evolution $\Delta\mu(t)$. The values of $I_{c}=135\,(10^{3}/\mathrm{s})$ and $\hbar G=59.8$ are determined from the simulated current-chemical potential relation [65]. With $I_{c}$ and the effective charging energy $E_{C}=4(\partial\mu/\partial N)$ we estimate the Josephson frequency $\omega_{\text{J}}=\sqrt{I_{c}E_{C}/\hbar}$ and obtain $f_{\text{J}}=38\,\mathrm{Hz}$ in agreement with the value of $f_{0}$. The result of the spectrum $|\Delta\mu(\omega)|^{2}$ is shown in Fig. 3d, which captures the dynamical response of the JPA. We note that the circuit model also describes higher JPA resonance and other excitation peaks well [65].
Figure 5: Dephasing dynamics. (a) Average vortex number $N_{v}$ as a function of $A_{\mathrm{d}}$ for the parameters used in Fig. 4(a). (b, c) Time evolution of the junction phase $\phi(t)$ for two different samples at $\tilde{I}_{1}=0.42$ with (continuous line) and without pump (dashed line). (d) Imbalance $z(t)$, averaged over many samples, corresponding to the parameters in (b, c).
To quantify the overall amplification, we define a dimensionless gain $G_{\mathrm{s}}=S_{\omega}(A_{\mathrm{d}})/S_{\omega}(0)-1$, where $S_{\omega}(A_{\mathrm{d}})$ represents the signal peak value and $S_{\omega}(0)$ corresponds to the result without pump. In Fig. 4(a) we demonstrate that $G_{\mathrm{s}}$ initially increases approximately quadratically with $A_{\mathrm{d}}$, and then reaches a maximum for a value of $A_{\mathrm{d}}$, that decreases with increasing $\tilde{I}_{1}=I_{1}/I_{c}$. For larger values of $A_{\mathrm{d}}$, the gain $G_{\mathrm{s}}$ decreases due to the dephasing dynamics as explained in the subsequent discussion. While the nonlinear increase is captured by the circuit model, the model fails to describe the decrease beyond the maximum, primarily visible for larger $\tilde{I}_{1}$. Fig. 4(b) presents the signal-to-noise ratio (SNR), defined as $\bar{S}_{\omega}/\Delta\bar{S}_{\omega}$, where we compute the mean signal value $\bar{S}_{\omega}=\left\langle S_{\omega}(A_{\mathrm{d}})-S_{\omega}(0)\right\rangle$ and the noise $\Delta\bar{S}_{\omega}$ from the variance determined by $\Delta\bar{S}_{\omega}^{2}=\left\langle(S_{\omega}(A_{\mathrm{d}})-S_{\omega}(0)-\bar{S}_{\omega})^{2}\right\rangle$. Here, $\langle..\rangle$ denotes a statistical average over $100$ samples of the ensemble. The SNR depends on both $A_{\mathrm{d}}$ and $\tilde{I}_{1}$, with a visible increase with increasing $A_{\mathrm{d}}$. The maximal regime of the SNR is in the same regime as the maximal gain $G_{\mathrm{s}}$, suggesting that this is the optimal regime of operation. In Fig. 4(c, d), the simulation results obtained at a higher temperature of $T/T_{0}=0.3$ show a noticeable reduction in overall gain, while still exhibiting a qualitatively similar trend to that observed at $T/T_{0}=0.06$. Thermal fluctuations suppress the Josephson plasma frequency $f_{0}$ and the critical current $I_{c}$, resulting in $f_{0}=35.2\,\mathrm{Hz}$ and $I_{c}=103\,(10^{3}/\mathrm{s})$ for $T/T_{0}=0.3$, see End Matter and [65] for details.
Figure 6: Underlying nonlinear mechanism. (a) Time evolution of the density $\delta n(x,t)=n({\bf r},t)-n({\bf r},0)$, averaged over the $y$ direction and the ensemble, is shown for $\tilde{I}_{1}=0.21$ without the pump. (b) Dynamics in the presence of pump with $A_{\mathrm{d}}=0.8$. (c) Ensemble averaged phase difference $\phi$ at the junction for the cases in (a) and (b). (d) The results of zeroth and first order expansion of Eq. 4 without the resistive term; see text.
Fig. 5(a) shows the average total number of vortices in the system, determined by calculating the local phase winding around each lattice plaquette and averaging over both the time evolution and the initial ensemble, see End Matter for details. The increase in vortex number coincides with the slowdown and eventual reduction of the gain observed in Fig. 4(a), confirming that vortex excitations are the primary dissipation mechanism limiting the amplification. During time evolution, the junction phase undergoes phase-slippage events, where the local phase difference at the barrier exceeds $\pi$, see Figs. 5(b, c). These events result in the creation of vortex-antivortex pairs near the junction, providing a dissipation channel for the injected pump energy. When averaged over many realizations, these stochastic slippage events give rise to dephasing in the imbalance dynamics, as shown in Fig. 5(c). This dephasing reflects the cumulative impact of vortex formation, which gradually suppresses coherent amplification and ultimately sets the limit for the achievable gain.
To understand the underlying nonlinear mechanism of the JPA we examine the time evolution of the density and the phase at the junction in Fig. 6. Without pump, the barrier velocity modulation results in a smooth excitation of the density, which is created by the change of junction phase that is synchronized with the signal frequency. When the pump field is on, there are higher-order density excitations that result from the modulation of the phase by the pump field. These can be explained with a perturbative solution of the equation $\tilde{I}_{1}\cos(\omega_{\mathrm{s}}t)=\bigl(1+A_{\mathrm{d}}\sin(\omega_{\mathrm{d}}t)\bigr)\sin\phi$, assuming a small modulation parameter $A_{\mathrm{d}}$ and expanding $\phi$ as $\phi=\phi_{0}+A_{\mathrm{d}}\phi_{1}+\mathcal{O}(A_{\mathrm{d}}^{2})$. At zeroth order we obtain $\phi_{0}\approx\tilde{I}_{1}\cos(\omega_{\mathrm{s}}t)$, which explains the smooth oscillation of the phase without the pump. The first-order term gives $\phi_{1}\approx-\tilde{I}_{1}\cos(\omega_{\mathrm{s}}t)\sin(\omega_{\mathrm{d}}t)\bigr)$, highlighting the intrinsic nonlinear mixing between the pump and signal field. In Fig. 6(d), the perturbative solutions consistently capture the dynamics of our atomic JPA.
Conclusions and outlook.— We have studied the dynamics of a driven atomic Josephson junction (JJ) using classical-field simulations. The JJ is formed by coupling two two-dimensional atom clouds with a tunnel barrier, where we periodically modulate both the position (input signal) and the height (pump) of the barrier. By tuning the pump frequency to twice the Josephson plasma frequency, we have demonstrated parametric amplification for the signal. Our system works in the $\mathrm{Hz}$ frequency range, that is beyond the operating range of the superconducting JPAs, thereby offering potential for low-frequency signal detection and precision sensing of slowly varying fields, such as gravitational or magnetic fields [Sheng2017, Javor2021, Somiya2023]. The current cold-atoms technology allows to accurately image the microscopic dynamics of the system [57, 58]. In this case, we have an unprecedented access to the density and phase microscopic dynamics at the junction responsible for the amplification and the intrinsic nonlinear mixing between the signal and pump fields. Our results can be experimentally probed using in-situ imaging and matter-wave interferometry, where the former enables measurement of the density imbalance and the latter allows detection of the underlying JPA mechanism. Our proposed protocol can be implemented in driven Josephson junctions experimentally realized in [57, 58].
Our findings have broad implications for the understanding of the basic mechanism on which parametric amplification relies in quantum systems. For the specific case of quantum electronics devices, while the microscopic phase-density dynamics can be inferred through suitable Josephson current measurements, their direct experimental probing is out of the reach of current techniques, such as scanning probe microscopy.
Acknowledgments.— L. M. acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG) in the excellence cluster ‘Advanced Imaging of Matter’ - EXC 2056 - project ID 390715994 and by ERDF of the European Union and ‘Fonds of the Hamburg Ministry of Science, Research, Equalities and Districts (BWFGB)’.